PART A: An electric current flows through a resistor. Which side will have the higher electric potential?

Select the correct answer

The - side (i.e. where the current technically flows out).

The + side (i.e. where the current technically flows in).

Both sides are exactly at the same potential.

PART B: If you connect two 100 W light bulbs in series to one 120 V outlet, what will be the total dissipated power (in both bulbs)?

Select the correct answer

100 W

less than 100 W

200 W

more than 200 W

PLEASE EXPLAIN THE CONCEPTS.
Answer #1

a)

the - side is the correct option

=====

b)

resistance of each bulb

R = 120^2 / 100 = 144 ohm

net resistance of circuit

Rnet = 288 ohm

current in circuit

i = V/Rnet = 120 / 288 = 0.417 A

power dissipated in each bulb

P = i^2 R = 0.417^2* 144 = 25 W

net power

Pnet = 2 P = 50 W

so less than 100W

=====

comment before rate in case any doubt, will reply for sure.. goodluck